# Multilayered Intelligent Reflecting Surface for Long-Range Underwater Acoustic Communication

### v1

@@ VII Conclusion
In this paper, a multilayered reconfigurable intelligent surfaces is proposed for underwater acoustic networks. The stacked architecture enables each individual reflector to generate a reflected wave with flexible amplitude and phase via IQ modulation. Compared with the multi-bit coding RIS commonly employed in terrestrial communications, the proposed ML-ARIS achieves higher-precision beam steering while minimizing interference to the surrounding environment. To ensure efficient operation in complex ocean settings, an enhanced matching network is also developed to maintain a stable output impedance under varying water temperatures and RIS deployment depths.
@@
The performance of the enhanced matching circuit and the multilayered acoustic reflector is evaluated through both tank experiments and COMSOL simulations around 28 kHz and 41 kHz. The results verify the effectiveness of the ML-ARIS in manipulating the phase and amplitude of the reflected wave to support advanced array processing algorithms. This capability positions ML-ARIS as a promising technology to improve the communication range and data rate for future underwater acoustic networks.

### v2

@@ VII Conclusion and Future Work
In this paper, a multilayered reconfigurable intelligent surfaces is proposed for underwater acoustic networks. The stacked architecture enables each individual reflector to generate a reflected wave with flexible amplitude and phase via synthetic reflection. Compared with the multi-bit coding RIS commonly employed in terrestrial communications, the proposed ML-ARIS achieves higher-precision beam steering while minimizing interference to the surrounding environment. To ensure efficient operation in complex ocean settings, an enhanced matching network is also developed to maintain a stable output impedance under varying water temperatures and RIS deployment depths.
@@
The performance of the enhanced matching circuit and the multilayered acoustic reflector is evaluated through both tank experiments and COMSOL simulations around 28 kHz and 41 kHz. The results verify the effectiveness of the ML-ARIS in manipulating the phase and amplitude of the reflected wave to support advanced array processing algorithms. This capability establishes ML-ARIS as an enabling hardware platform for future studies of RIS-assisted underwater acoustic communication systems.
UA-RIS is still an emerging research area, and many of the supporting elements required for fielded RIS-assisted links remain open. While this work focuses on the proposed UA-RIS architecture and its hardware foundations, enabling high-resolution reflection control and improving matching robustness under underwater impedance variability, we emphasize that realizing end-to-end communication gains in practical waveguide environments will require further system-level development. In particular, future work should target supporting protocols and signal processing, including channel sounding and tracking, pilot and feedback design, configuration update strategies for time-varying environments and mobility, and practical coordination mechanisms between transmitter, UA-RIS, and receiver to enable reliable RIS-assisted underwater acoustic communications in operational deployments.
Acknowledgement
This work is supported in part by the US National Science Foundation under Awards CNS-2048188 and CNS-2016726 and by the Alabama Water Institute Equipment Fund, The University of Alabama.